# Robust Preference Optimization through Reward Model Distillation

### v1

@@ Abstract
Language model (LM) post-training (or alignment) involves maximizing a reward function that is derived from preference annotations. Direct Preference Optimization (DPO) is a popular offline alignment method that trains a policy directly on preference data without the need to train a reward model or apply reinforcement learning. However, typical preference datasets have only a single, or at most a few, annotation per preference pair, which causes DPO to overconfidently assign rewards that trend towards infinite magnitude. This frequently leads to degenerate policies, sometimes causing even the probabilities of the preferred generations to go to zero. In this work, we analyze this phenomenon and propose distillation to get a better proxy for the true preference distribution over generation pairs: we train the LM to produce probabilities that match the distribution induced by a reward model trained on the preference data. Moreover, to account for uncertainty in the reward model we are distilling from, we optimize against a family of reward models that, as a whole, is likely to include at least one reasonable proxy for the preference distribution. Our results show that distilling from such a family of reward models leads to improved robustness to distribution shift in preference annotations, while preserving the simple supervised nature of DPO.
11footnotetext: Core contributor. Correspondence to fisch@google.com.
@@ 1 Introduction
Language model (LM) post-training (or alignment) aims to steer language model policies towards responses that agree with human preferences. Early state-of-the-art approaches have focused on reward learning from human feedback. In this paradigm, preference annotations are used to train reward models, which then guide the optimization of the language model policy through online reinforcement learning (an approach broadly referred to as RLHF). Recent research on offline “Direct Preference Optimization” [DPO; 23] and extensions thereof [3; 31], however, has demonstrated that it is also possible to directly optimize policies on the preference data, which bypasses the need for a separate reward model—and its offline nature also leads to faster, and simpler, training frameworks.
@@
While this direct approach to preference optimization is attractive in terms of its simplicity and efficiency, it also raises important questions about the effectiveness and robustness of the resulting policies—as well as the broader utility of using an explicit reward model. In this paper, we argue that explicit reward modeling can, in fact, offer substantial practical advantages that are not captured by DPO’s formulation. In particular, we theoretically show that relying solely on the preference data can be a precarious strategy, with few natural brakes in place to prevent policies trained under the DPO objective from careening off towards degenerate policies when the preference data exhibits certain idiosyncratic properties. On the other hand, explicit reward models can easily be regularized and understood—regardless of whether they are Bradley-Terry models [4], margin-based ranking models [40], or simply any other kind of function that correlates well with human preferences [31; 17].
@@
Taking a step back from pure direct preference optimization, we propose a method that merges the best of both worlds: an efficient reward model distillation algorithm that (i) operates effectively in the offline setting, (ii) makes minimal assumptions about the true, optimal reward we aim to maximize, and (iii) demonstrates greater robustness to the specific distribution of prompt/response data used for policy alignment. Drawing inspiration from prior knowledge distillation techniques [14; 26; 35; 10], we leverage the same change of variables trick employed in DPO to express the language model policy in terms of its implicit reward model [23]. We then train the policy to match our desired, explicit reward via an $L_{2}$ loss that directly regresses the pairwise differences in target rewards for any two generation pairs $(x,y_{1})$ and $(x,y_{2})$. We theoretically establish the equivalence between optimizing this simple distillation loss over a sufficiently diverse offline dataset of unlabeled examples and optimizing the traditional online RLHF objective with reinforcement learning.
@@
Our reward model distillation approach, however, is not immune to some of the same challenges facing DPO-style learning of policies. In particular, reward model distillation requires having a reliable reward model—but having a reliable reward requires having a reliable method for extracting a reward model from a potentially noisy preference dataset. To address the uncertainty surrounding the “right” reward model, we introduce a pessimistic extension to our approach. This extension aims to maximize the worst-case improvement of our model across a plausible family of reward models (e.g., those sufficiently consistent with annotated preference data). This strategy aligns with that of existing work in conservative offline reinforcement learning [5; 16]. Interestingly, we derive that this pessimistic objective can be equivalently expressed and optimized by adding a simple additional KL-divergence regularization to the original distillation objective.
@@
Empirically, we find that reward model distillation, particularly pessimistic reward model distillation, leads to similar performance to prior direct preference optimization methods in settings where the preference datasets used are unbiased, but significantly better performance in settings where the preference datasets are biased, when compared to DPO and the Identity Preference Optimization (IPO) framework of [3], which was introduced as a more robust alternative to DPO. To further support these empirical observations, we provide an extensive theoretical analysis that both (i) sheds more light on the degenerative tendencies of DPO and issues inherent to its objective, and (ii) highlights relative advantages of our explicitly regularized approaches.
@@ 2 Preliminaries
We begin with a brief review of Direct Preference Optimization (DPO) [23] and its analysis. Proofs of all theoretical results provided here, and in the rest of the paper, are deferred to Appendix A.
@@ 2.1 The preference alignment problem
Let $x$ be an input prompt, and let $y\sim\pi_{\theta}(\cdot\mid x)$ be the language model policy $\pi_{\theta}$’s response to $x$. Given some reward function $r^{*}(x,y)$ and another reference policy $\pi_{\mathrm{ref}}(y\mid x)$, the goal of alignment is to solve for the “aligned” policy $\pi_{\theta^{*}}(y\mid x)$ that maximizes the following RLHF objective, i.e.,
@@
where $\mu(x)$ is a fixed distribution over prompts, and the KL-divergence term prevents the aligned policy from being dramatically different from the anchoring reference policy, $\pi_{\mathrm{ref}}(y\mid x)$. Here, the reward function $r^{*}$ is typically not known in advance, but rather inferred from collected human preference data in the form of $(x,y^{w},y^{\ell})$, where $x$ is the prompt, $y^{w}$ is the “winning”, or preferred, response, and $y^{\ell}$ is the “losing”, or dispreferred, response. A common approach is to assume that pairs $(y_{1},y_{2})$ follow a Bradley-Terry model [4], under which the probability that $y_{1}$ is preferred to $y_{2}$ given the reward function $r^{*}$ and prompt $x$ is $p^{*}(y_{1}\succ y_{2}\mid x)=\sigma(r^{*}(x,y_{1})-r^{*}(x,y_{2}))$, where $\sigma(\cdot)$ is the sigmoid function and $\succ$ denotes preference. Under this model, we can use the preference data $(x,y^{w},y^{\ell})\sim\mathcal{D}_{\mathrm{pref}}$ to estimate $r^{*}$ via maximum likelihood estimation, i.e.,
$\hat{r}\in\arg\!\min_{r}\mathbb{E}_{(y^{w},y^{\ell},x)\sim\mathcal{D}_{\mathrm{pref}}}\left[-\log\sigma(r(x,y^{w})-r_{\phi}(x,y^{\ell}))\right].$ (2)
With $\hat{r}$ in hand, Eq. (1) can be optimized using standard reinforcement learning algorithms [27; 29; 6].
@@ 2.2 Direct preference optimization
DPO is a simple approach for offline policy optimization that uses preferences to directly align the language model policy, without training an intermediate reward model. Specifically, DPO leverages the fact that the optimal solution to the KL-constrained objective in (1) takes the form [15]
@@
$p^{*}(y_{1}\succ y_{2}\mid x)=\sigma\left(\beta\log\frac{\pi_{\theta^{*}}(y_{1})\pi_{\mathrm{ref}}(y_{2})}{\pi_{\theta^{*}}(y_{2})\pi_{\mathrm{ref}}(y_{1})}\right),$ (5)
@@
$\mathcal{L}_{\mathrm{dpo}}(\pi_{\theta};\mathcal{D}_{\mathrm{pref}})=\mathbb{E}_{(y^{w},y^{\ell},x)\sim\mathcal{D}_{\mathrm{pref}}}\left[-\log\sigma\left(\beta\log\frac{\pi_{\theta^{*}}(y^{w})\pi_{\mathrm{ref}}(y^{\ell})}{\pi_{\theta^{*}}(y^{\ell})\pi_{\mathrm{ref}}(y^{w})}\right)\right].$ (6)
2.3 Pitfalls of direct preference optimization
As argued in [3], the Bradley-Terry assumption that DPO strongly relies on for maximum likelihood estimation is sensitive to the underlying preference data. Specifically, if we have any two responses $y_{1}$ and $y_{2}$ where $p^{*}(y_{1}\succ y_{2}\mid x)=1$, then the Bradley-Terry model dictates that $r^{*}(y_{1})-r^{*}(y_{2})=+\infty$, and therefore $\pi_{\theta^{*}}(y_{2}\mid x)=0$ for any finite KL-regularization strength $\beta$.
We can illustrate this phenomenon on a broader level with the following example.
@@ 2.3 Pitfalls of direct preference optimization
Additional analysis of the training dynamics of DPO is also provided in §5. A significant, and non-obvious, implication of Corollary 1 is that the set of global optima of the DPO loss also includes policies that can shift nearly all probability mass to responses that never even appear in the training set—and even assign near zero probability to all of the training data responses that do in fact correspond to winning generations, $y^{w}$, a phenomenon that has been observed empirically [e.g., 20]. Stated differently, Corollary 1 implies that any $\theta^{*}$ merely satisfying $\pi_{\theta^{*}}(y^{\ell}_{i})=0$ with $\pi_{\theta^{*}}(y^{w}_{i})>0$ $\forall i\in[n]$ is a global minimizer of the DPO objective in this setting. Though simplistic, the scenario in Assumption 1 is closer to reality than might first be appreciated: in many practical situations we can almost always expect the finite-sample preference data to contain one (or at most a few) preference annotations per example $(x,y_{1},y_{2})$, while the policies $\pi_{\theta}$ can have billions of parameters ($\gg n$). Of course, this issue can also be viewed as a classic instance of overfitting—with the additional caveat that as opposed to overpredicting responses within the training set, we might overfit to almost never producing anything like the “good” responses that do appear within the training set. Furthermore, without additional regularization (beyond $\beta$), we can expect this degeneration to easily happen in typical preference datasets.

### v2

@@ Abstract
Language model (LM) post-training (or alignment) involves maximizing a reward function that is derived from preference annotations. Direct Preference Optimization (DPO) is a popular offline alignment method that trains a policy directly on preference data without the need to train a reward model or apply reinforcement learning. However, the empirical evidence suggests that DPO typically assigns implicit rewards that overfit, and trend towards infinite magnitude. This frequently leads to degenerate policies, sometimes causing even the probabilities of the preferred generations to go to zero. In this work, we analyze this phenomenon and use distillation to get a better proxy for the true preference distribution over generation pairs: we train the LM such that its induced implicit reward, i.e., the scaled log-likelihood ratio of the model to the reference model, matches an explicit reward model trained on the preference data. Moreover, to account for uncertainty in the reward model we are distilling from, we optimize against a family of reward models that, as a whole, is likely to include at least one reasonable proxy for the preference distribution. Our results show that distilling from such a family of reward models leads to improved robustness to distribution shift in preference annotations, while preserving the simple supervised nature of DPO.
11footnotetext: Core contributor.
@@ 1 Introduction
Language model (LM) post-training (or alignment) aims to steer language model policies towards responses that agree with human preferences. Early state-of-the-art approaches have focused on reward learning from human feedback. In this paradigm, preference annotations are used to train reward models, which then guide the optimization of the language model policy through online reinforcement learning (an approach broadly referred to as RLHF). Recent research on offline “Direct Preference Optimization” (DPO; Rafailov et al., 2023) and extensions thereof Azar et al. (2024); Tang et al. (2024b); Meng et al. (2024), however, has demonstrated that it is possible to directly optimize policies on the preference data, which (a) bypasses the need for a separate reward model, and (b) uses standard supervised techniques rather than online reinforcement learning, which can be more difficult to optimize. These advantages have led to the the adoption of offline alignment, and in particular offline DPO, as the post-training algorithm of choice in both smaller-scale academic settings as well as larger-scale projects such as Llama 3 AI@Meta (2024) and OLMo Groeneveld et al. (2024).
@@
While this direct approach to preference optimization is attractive in its simplicity and efficiency, it also raises questions about the effectiveness and robustness of the resulting policies—as well as the broader utility of an explicit reward model beyond online reinforcement learning. In this paper, we argue that explicit reward modeling can, in fact, offer substantial practical and theoretical benefits. In particular, we theoretically show that relying solely on the preference data can be a precarious strategy, with few natural brakes in place to prevent policies trained under the DPO objective from careening off towards degenerate policies when the preference data exhibits certain idiosyncratic properties. On the other hand, explicit reward models can easily be regularized and understood—regardless of whether they are Bradley-Terry models Bradley and Terry (1952), margin-based ranking models Zhao et al. (2023), or any other function that correlates well with human preferences Lee et al. (2023); Tang et al. (2024b); Swamy et al. (2024).
@@
Taking a step back from pure direct preference optimization, we first explore a method that merges the best of both worlds for the offline setting: an efficient reward model distillation algorithm that (i) operates effectively in the offline setting, (ii) makes minimal assumptions about the true, optimal reward we aim to maximize, and (iii) demonstrates greater robustness to the specific distribution of prompt/response data used for policy alignment. Drawing inspiration from prior knowledge distillation techniques Hinton et al. (2015); Romero et al. (2015); Yang et al. (2019); Furlanello et al. (2018), we use the same change of variables trick employed in DPO to express the language model policy in terms of its implicit reward model Rafailov et al. (2023). We then train the policy’s implicit reward model to match our desired, explicit reward via an $L_{2}$ loss that directly regresses the pairwise differences in target rewards for any two generation pairs $(x,y_{1})$ and $(x,y_{2})$. [^1] We also note that a similar $L_{2}$ reward alignment loss was recently explored in independent work by Mao et al. (2024) and Gao et al. (2024b), giving further support for its effectiveness. See a discussion of related work in §2. Here we theoretically establish the equivalence between optimizing this simple distillation loss over a sufficiently diverse offline dataset of unlabeled examples, and optimizing the traditional online RLHF objective with reinforcement learning.
While this approach adds reward modeling and reward inference back into the pipeline, it still maintains much of the simplicity and efficiency of reward-model-free DPO. Specifically, in our setting, rewards for the training data can be computed offline, once, ahead of time. This computation is completely parallelizable, and reward inference is significantly faster than the autoregressive generation (also known as model rollout) that is done in online settings. Consequently, this allows the policy training framework to be nearly identical to standard DPO, modulo the structure of the data that is fed in. This is true regardless of the size of the policy that is trained (as the policy’s implicit reward model does not need to be the same size as the explicit reward model) or its hyper-parameters (as the policy’s hyper-parameters do not need to match those of the reward model).
@@
Reward model distillation, however, is still subject to some of the same challenges facing DPO-style training. In particular, distillation requires having a reliable reward model—but having a reliable reward still requires having a reliable method for extracting a reward model from a potentially noisy preference dataset. To address the uncertainty surrounding what the “right” reward model to optimize against is, we also introduce a pessimistic extension to our approach. This extension aims to maximize the worst-case improvement of our model across a plausible family of reward models (e.g., those sufficiently consistent with annotated preference data). This strategy aligns with that of existing work in conservative offline reinforcement learning Cheng et al. (2022); Kumar et al. (2020). We show that this pessimistic objective can be equivalently expressed and optimized by adding a simple additional KL-divergence regularization to the original distillation objective.
@@
Empirically, we find that reward model distillation, particularly pessimistic reward model distillation, leads to similar performance to prior direct preference optimization methods when the preference datasets used are unbiased. When the preference datasets are biased, however, it leads to significantly better performance when compared to DPO and the Identity Preference Optimization (IPO) framework of Azar et al. (2024), which was introduced as a more robust alternative to DPO. To further support these empirical observations, we provide an extensive theoretical analysis that both (i) sheds more light on the degenerative tendencies of DPO and issues inherent to its objective, and (ii) highlights relative advantages of our explicitly regularized approaches.
2 Related work
Recent work in offline alignment has focused on DPO Rafailov et al. (2023) as a simpler alternative for aligning language models from preference data. Subsequent work, however, has identified issues with DPO, including weak regularization Azar et al. (2024) and a tendency to decrease the probability of winning generations during training Pal et al. (2024). Similar to some of the findings in this work, Rafailov et al. (2024a) theoretically showed the existence of minima to the empirical DPO loss that assign non-zero probability to outputs not present in the training data, which can lead to unexpected model behaviour. Here we further show that not only can such minima place non-zero probability on outputs not present in the training data, but they can also assign near-zero probability to preferred responses that do appear in the training data. A number of methods have since explored various avenues for combating these issues. These include analyzing the impact of noise on DPO alignment Gao et al. (2024a), proposing to update the reference policy during training Gorbatovski et al. (2024), and suggesting a variant of IPO with a per-context margin Amini et al. (2024). Additional research has focused on token-level alignment methods Zeng et al. (2024); Rafailov et al. (2024b); Mudgal et al. (2024); Chakraborty et al. (2024) and on developing a unified view of various offline alignment methods Tang et al. (2024b). Similar to our contributions towards using pessimism with respect to the correct choice of reward model for robust reward model distillation, existing work on offline RLHF has also focused on encompassing various forms of conservative reward penalties Zhu et al. (2023); Liu et al. (2024); Zhan et al. (2024). Most relevant to our work on a technical level, concurrent work by Mao et al. (2024) also uses an offline loss that leverages targets from a learned reward model for supervision (as opposed to binary preference labels) that has the same form of the simple “distillation loss” analyzed in §5.1 of this paper, but does not explore pessimism. Similarly, the REBEL algorithm of Gao et al. (2024b) also studies the same reward regression loss, but in the online setting. This work builds upon these findings, and provides further analysis, as well as a solution based on pessimism together with reward distillation.
As discussed in §1, offline settings are attractive since they allow for simple, efficient, and scalable training frameworks. At the same time, while offline alignment methods are popular, recent evidence suggests that online alignment methods such as RLHF Christiano et al. (2017); Stiennon et al. (2020), can still lead to more favorable outcomes Guo et al. (2024); Tajwar et al. (2024); Dong et al. (2024); Xu et al. (2024); Xiong et al. (2024); Calandriello et al. (2024), especially when high reward outputs have low probability under the base policy. One of the advantages of online settings over offline settings is that many strategies for mitigating over-optimization that are simple to apply in online settings, such as reward shaping, are not as straightforward to apply in offline settings. For example, reward ensembles have been widely investigated recently as a mechanism for tackling reward hacking in online RLHF Eisenstein et al. (2023); Coste et al. (2023); Zhai et al. (2023); Ramé et al. (2024), and in the context of multi-objective optimization Moskovitz et al. (2023); Rame et al. (2024). Other notable examples for combating reward over-fitting and optimization by training reward models with regularized training objectives include iterative data smoothing Zhu et al. (2024), which uses a trained model to softly label data during RLHF, and reward calibration from demonstrations Rita et al. (2024). This work addresses some of the methodological and experimental gap that exists between online and offline methods for RLHF, by allowing for explicitly designed, trained, and regularized reward models (or pessimistic reward model ensembles) to be added back into the offline alignment setting without losing the practical benefits of the offline setting. Also relevant to our work, Moskovitz et al. (2023) focus on “composite” rewards in the online setting, with the goal of achieving high task reward while ensuring that every individual component is above some threshold—also by applying a Lagrangian relaxation. In this work, we also consider multiple reward models, but we only focus on cases where there is no known, obvious reward decomposition.
Finally, the question of using a small amount of offline data to learn high-quality policies, instead of online access to reward feedback, has also been widely studied in the offline reinforcement learning (RL) literature. The predominant approach here is to use pessimism, that is, to learn a policy with the highest reward under all plausible environment models consistent with the data, with an extensive theoretical (Liu et al., 2020; Zanette et al., 2021; Xie et al., 2021) and empirical (Kumar et al., 2020; Cheng et al., 2022; Yu et al., 2021) body of supporting work. The key insight in this literature is that without pessimism, the RL algorithm learns undesirable behaviors which are not explicitly ruled out in the training data, and pessimism provides a robust way of preventing such undesirable extrapolations, while still preserving generalization within the support of the data.
@@ 3 Preliminaries
We begin with a brief review of Direct Preference Optimization (DPO) Rafailov et al. (2023).
@@ 3.1 The preference alignment problem
Let $x$ be an input prompt, $y\sim\pi_{\theta}(\cdot\mid x)$ be the language model policy $\pi_{\theta}$’s response to $x$, and $\pi_{\mathrm{ref}}(y\mid x)$ a reference policy (such as a pretrained or finetuned language model that is high-performing, but not yet aligned, and often used as the starting point for optimization). Given some reward function $r^{*}(x,y)$, the goal of alignment is to solve for the “aligned” policy $\pi_{\theta^{*}}(y\mid x)$ that maximizes the following RLHF objective, i.e.,
@@
where $\mu(x)$ is a fixed distribution over prompts, and the KL-divergence term keeps the aligned policy close to the anchoring reference policy, $\pi_{\mathrm{ref}}(y\mid x)$. Here, the reward function $r^{*}$ is typically not known in advance, but rather inferred from collected human preference data in the form of $(x,y^{w},y^{\ell})$, where $x$ is the prompt, $y^{w}$ is the “winning”, or preferred, response, and $y^{\ell}$ is the “losing”, or dispreferred, response. A common approach is to assume that $(y_{1},y_{2})$ follow a Bradley-Terry model Bradley and Terry (1952), under which the probability that $y_{1}$ is preferred to $y_{2}$ given the reward function $r^{*}$ and prompt $x$ is $p^{*}(y_{1}\succ y_{2}\mid x)=\sigma(r^{*}(x,y_{1})-r^{*}(x,y_{2}))$, where $\sigma(\cdot)$ is the sigmoid function and $\succ$ denotes preference. Under this model, we can use the preference data $(x,y^{w},y^{\ell})\sim\mathcal{D}_{\mathrm{pref}}$ to estimate $r^{*}$ via maximum likelihood estimation, i.e.,
$\hat{r}\in\arg\!\min_{r}\mathbb{E}_{(y^{w},y^{\ell},x)\sim\mathcal{D}_{\mathrm{pref}}}\left[-\log\sigma(r(x,y^{w})-r(x,y^{\ell}))\right].$ (2)
With $\hat{r}$ in hand, Eq. (1) can be optimized using standard reinforcement learning algorithms Schulman et al. (2017); Stiennon et al. (2020); Christiano et al. (2017).
@@ 3.2 Direct preference optimization
DPO is a simple approach for offline policy optimization that uses preferences to directly align the language model policy, without training an intermediate reward model. Specifically, DPO leverages the fact that the optimal solution to the KL-constrained objective in (1) takes the form Korbak et al. (2022)
@@
$p^{*}(y_{1}\succ y_{2}\mid x)=\sigma\left(\beta\log\frac{\pi_{\theta^{*}}(y_{1}\mid x)\pi_{\mathrm{ref}}(y_{2}\mid x)}{\pi_{\theta^{*}}(y_{2}\mid x)\pi_{\mathrm{ref}}(y_{1}\mid x)}\right),$ (5)
@@
$\mathcal{L}_{\mathrm{dpo}}(\pi_{\theta};\mathcal{D}_{\mathrm{pref}})=\mathbb{E}_{(y^{w},y^{\ell},x)\sim\mathcal{D}_{\mathrm{pref}}}\left[-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y^{w}\mid x)\pi_{\mathrm{ref}}(y^{\ell}\mid x)}{\pi_{\theta}(y^{\ell}\mid x)\pi_{\mathrm{ref}}(y^{w}\mid x)}\right)\right].$ (6)
As described in §1, optimizing $\mathcal{L}_{\mathrm{dpo}}$ offers two main advantages over using online RL for Eq. (1): (a) there is no need for a separate reward model, and (b) $\mathcal{L}_{\mathrm{dpo}}$ is a supervised objective that can be trained offline, which allows for a simpler training setup than online learning. Still, $\mathcal{L}_{\mathrm{dpo}}$ also has certain pitfalls, as we analyze next.
4 Pitfalls of direct preference optimization
As argued by Azar et al. (2024), DPO strongly relies on the Bradley-Terry assumption, which leads to surprising and undesirable consequences when trained on finite preference data. The root issue is that if we have any two responses $y_{1}$ and $y_{2}$ where $p^{*}(y_{1}\succ y_{2}\mid x)=1$, then the Bradley-Terry model dictates that $r^{*}(y_{1})-r^{*}(y_{2})=+\infty$, and therefore $\pi_{\theta^{*}}(y_{2}\mid x)=0$ for any finite KL-regularization strength $\beta$.
We can illustrate this phenomenon on a broader level with the following example:
@@ 4 Pitfalls of direct preference optimization
Additional analysis of the training dynamics of DPO is provided in §7. A significant implication of this result is that the set of global optima of the DPO loss includes policies that can shift nearly all probability mass to responses that never appear in the training set—and even assign near-zero probability to all of the training data responses that do in fact correspond to winning generations, $y^{w}$, a phenomenon that has been observed empirically and analyzed theoretically Pal et al. (2024); Rafailov et al. (2024a; b); Tajwar et al. (2024). [^2] In fact, the same conclusions can also be inferred from Proposition 1 in the concurrent work of Rafailov et al. (2024a). While their assumptions are slightly different, similarly to us, they also target scenarios in which the training data exhibits certain (realistic) deficiencies.
@@
Stated differently, Corollary 1 implies that any $\theta^{*}$ merely satisfying $\pi_{\theta^{*}}(y^{\ell}_{i})=0$ with $\pi_{\theta^{*}}(y^{w}_{i})>0$ $\forall i\in[n]$ is a global minimizer of the DPO objective in this setting. Though simplistic, the scenario in Assumption 1 is closer to reality than might first be appreciated: in many practical situations we can expect the finite-sample preference data to contain one (or at most a few) preference annotations per example $(x,y_{1},y_{2})$, while the policies $\pi_{\theta}$ can have billions of parameters ($\gg n$). It is important to note that the supposed regularization term $\beta$ does not help: it can limit the speed at which the optimizer reaches the degenerate solution, but it cannot alter the final destination. This issue could be viewed as a classic instance of overfitting—but as opposed to overpredicting responses within the training set, we might overfit to almost never producing anything like the “good” responses that do appear within the training set. Furthermore, without additional regularization (beyond $\beta$), we can expect this degeneration to occur in typical preference datasets.
Takeaways: Pitfalls of direct preference optimization The DPO objective only requires that the likelihood of the preferred response is relatively higher than that of the dispreferred response. A peculiarity of this objective is that when preference data is disjoint (i.e., preferred responses never appear as dispreferred responses, and vice versa), certain types of policies (e.g., over-parameterized) will learn to assign 00 probability to all dispreferred responses, with merely non-zero probability to all preferred responses. This includes policies that assign near-zero probability to preferred responses, and place all mass on (often degenerate) generations outside the training set.
4.1 Case study: degenerate DPO optima in a BoW model
For intuition on why the DPO global optima can include policies where $\pi(y^{w})$ may be nearly 00 for all $y^{w}$ in the training set, consider the simplified case where the policy is a bag-of-words model, $\pi_{\theta}(y)\propto\exp\left(c(y)\cdot\theta\right)$ for $c(y)$ representing a vector of counts in $y$ and $\theta_{i}$ representing the unnormalized log-probability of token $i$. Then we can formally show that DPO optimization monotonically decreases an upper bound on the probability of the preferred completion, $\tilde{\pi}_{\theta^{(t-1)}}(y^{w})\geq\tilde{\pi}_{\theta^{(t)}}(y^{w})\geq\pi_{\theta^{(t)}}(y^{w}).$
Proposition 2.
Let $y^{w},y^{\ell}\in\mathcal{V}^{n}$ be preferred versus dispreferred outputs of length $n$, respectively, with $\pi_{\mathrm{ref}}(y^{w}),\pi_{\mathrm{ref}}(y^{\ell})>0$ and corresponding count vectors $c(y^{w}),c(y^{\ell}).$ Let $\log\pi_{\theta}(y)=c(y)\cdot\theta-nZ(\theta)$ for $Z(\theta)=\log\sum^{\mathcal{V}}_{i}e^{\theta_{i}}$, with upper bound $\log\tilde{\pi}_{\theta}(y)=c(y)\cdot\theta-n\max_{j}\theta_{j}$. Let $\theta^{(t)}$ represent the parameters of $\pi$ after $t$ steps of gradient descent on $\mathcal{L}_{\mathrm{dpo}}(\{y^{\ell},y^{w},x\})$, with $\theta^{(0)}=0$. Then, we have that $\pi_{\theta^{(t)}}(y^{w})\leq\tilde{\pi}_{\theta^{(t)}}(y^{w})\leq\tilde{\pi}_{\theta^{(t-1)}}(y^{w})$ for all $t$, with strict inequality when $||c(y^{w})-c(y^{\ell})||_{0}>1$.
If $\pi_{\theta^{(t)}}(y^{w})$ decreases in $t$, what other strings become more probable? In the following proposition, we show that under the bag-of-words model, DPO optimization moves probability mass away from $y^{w}$ to sequences that contain only the tokens that maximize the difference between $y^{w}$ and $y^{\ell}$.
Proposition 3.
Let $y^{w}$ and $y^{\ell}$ be preferred versus dispreferred outputs of length $n$. Let $\Delta=c(y^{w})-c(y^{\ell})$ be the difference in unigram counts. Let $\hat{y}=[i,i,\ldots,i]$, for $i\in\text{arg}\max\Delta,$ with $\|c(\hat{y})\|_{1}=n$. Then $\pi_{\theta^{(t)}}(y^{w})-\pi_{\theta^{(t)}}(\hat{y})=\tau(t)k$ for some $k\leq 0$ and some non-decreasing $\tau:\mathbb{Z}_{+}\to\mathbb{R}_{+}$.
We have $k=0$ when $c(y^{w})=c(\hat{y})$, and $k\ll 0$ when $\|c(y^{w})\|_{2}\ll\|c(\hat{y})\|_{2}=n$ (when $c(y^{w})$ is dense) and $\|\Delta\|_{2}\approx\|\Delta\|_{\infty}$ (when $\Delta$ is sparse). This implies that when $y^{w}$ and $y^{\ell}$ are similar, $\pi_{\theta}(y^{w})$ will degrade more rapidly. Early stopping will therefore have to trade off between reaching the degenerate solution on such cases, and underfitting other cases in which $y^{w}$ and $y^{\ell}$ are more distinct. Related findings are also reported in Razin et al. (2024).
Takeaways: BoW model case study For a simple BoW model, we show a realized setting in which DPO causes the probabilities of preferred outputs to catastrophically plummet, and instead puts all probability mass on generations that simply maximize token-level differences between preferred and dispreferred examples.